# Lower Bounds for Faithful Linear Representations of Subgroups of the Mapping Class Group

### v1

@@ 5. Lower Bounds for Faithful Representations of the Mapping Class Group
For each $e_{i}$, let $a_{i}^{\prime}\subseteq\Sigma_{g}$ be a meridian around $e_{i}$. These curves form a pants decomposition of $\Sigma_{g}$. By thickening the neighborhood $U_{i}$ of $e_{i}$ we obtain a neighborhood $S_{i}$ of $a_{i}^{\prime}$ which is a $4$-holed sphere. This neighborhood may also be obtained by gluing two adjacent pairs of pants along their common boundary component $a_{i}^{\prime}$. Let $b_{i}^{\prime}\subseteq\Sigma_{g}$ be as in Figure 8.
@@
(b) $a_{i}^{\prime},b_{i}^{\prime}\subseteq S_{i}$
Figure 8. Definition of the curves $a_{i}^{\prime}$ and $b_{i}^{\prime}$.
We can find representatives of $a_{i}^{\prime}$ and $b_{i}^{\prime}$ lying in $\Sigma\subseteq\Sigma_{g}$ and avoiding its marked points. We then take $a_{i},b_{i}\subseteq\Sigma$ as the isotopy classes of such representatives. Although such curves depend on a choice of representative of the curves $a_{i}^{\prime}$ and $b_{i}^{\prime}$, this choice is inconsequential to us. Since $\Gamma_{g}\setminus U_{i}$ is connected, so is $\Sigma_{g}\setminus S_{i}$. In particular, $a_{i}$ and $b_{i}$ are nonseparating.
@@
It is clear from Corollary 4.3 that $N_{j}M_{i}=M_{j}N_{i}=0$ for $i\neq j$ and $M_{j}M_{i}=0$ for all $i,j$. On the other hand, by assumption, $L_{a_{i}}$ and $L_{b_{i}}$ do not commute. Hence $M_{i}$ and $N_{i}$ don’t commute. In particular, $M_{i}N_{i}\neq 0$ or $N_{i}M_{i}\neq 0$. But $M_{i}N_{i}=0\iff N_{i}M_{i}=0$.
Indeed, the pairs $(T_{a_{i}},T_{b_{i}})$ and $(T_{b_{i}},T_{a_{i}})$ are conjugate in $\operatorname{PMod}(\Sigma)$: we can find $f\in\operatorname{PMod}(\Sigma)$ with $f(a_{i})=b_{i}$ and $f(b_{i})=a_{i}$, so that $fT_{a_{i}}f^{-1}=T_{b_{i}}$ and $fT_{b_{i}}f^{-1}=T_{a_{i}}$. Hence $M_{i}N_{i}$ and $N_{i}M_{i}$ are conjugated by $\rho(f)\in\operatorname{GL}_{d}(\mathbb{C})$. We are done. ∎
@@
range
pages10 rangepages3 rangepages1 rangepages52 rangepages72 rangepages6 rangepages44 rangepages12 rangepages1 rangepages19 rangepages17 rangepages20 rangepages14 rangepages36 rangepages18 rangepages7 rangepages14 rangepages37 rangepages10 rangepages1 rangepages9 rangepages47 rangepages30 rangepages34 rangepages4 rangepages53 rangepages37 rangepages33 rangepages18

### v2

@@ 5. Lower Bounds for Faithful Representations of the Mapping Class Group
For each $e_{i}$, let $a_{i}\subseteq\Sigma_{g}$ be a meridian around $e_{i}$. These curves form a pants decomposition of $\Sigma_{g}$. By thickening the neighborhood $U_{i}$ of $e_{i}$ we obtain a neighborhood $S_{i}$ of $a_{i}$ which is a $4$-holed sphere. This neighborhood may also be obtained by gluing two adjacent pairs of pants along their common boundary component $a_{i}$. Let $b_{i}\subseteq\Sigma_{g}$ be as in Figure 7.
@@
(b) $a_{i},b_{i}\subseteq S_{i}$
Figure 7. Definition of the curves $a_{i}$ and $b_{i}$.
We can find representatives of $a_{i}$ and $b_{i}$ lying on $\Sigma\subseteq\Sigma_{g}$ and avoiding its marked points, which we denote by $a_{i},b_{i}\subseteq\Sigma$ as well. Although such curves depend on a choice of representative of $a_{i},b_{i}\subseteq\Sigma_{g}$, this choice is inconsequential to us. Since $\Gamma_{g}\setminus U_{i}$ is connected, so is $\Sigma_{g}\setminus S_{i}$. In particular, $a_{i}$ and $b_{i}$ are nonseparating.
It is clear the curves $a_{i}$ and $a_{j}$ are disjoint for $i\neq j$. While $a_{i}$ is disjoint from $b_{j}$ for all $i\neq j$, a careful reader might notice the curves $b_{i}$ and $b_{j}$ can intersect each other nontrivially when $e_{i}$ and $e_{j}$ are adjacent edges of $\Gamma_{g}$. We will see this will not pose a problem to our proof.
@@
It is clear from Corollary 4.3 that $N_{j}M_{i}=M_{j}N_{i}=0$ for $i\neq j$ and $M_{j}M_{i}=0$ for all $i,j$. Since the curves $b_{i}$ and $b_{j}$ can intersect each other nontrivially, we cannot conclude $N_{i}N_{j}=0$. This fact is, however, immaterial to our proof, as the relations $N_{i}N_{j}=0$ are not part of the hypothesis of Proposition 3.8.
On the other hand, by assumption, $L_{a_{i}}$ and $L_{b_{i}}$ do not commute. Hence $M_{i}$ and $N_{i}$ don’t commute. In particular, $M_{i}N_{i}\neq 0$ or $N_{i}M_{i}\neq 0$. But $M_{i}N_{i}=0\iff N_{i}M_{i}=0$. Indeed, the pairs $(T_{a_{i}},T_{b_{i}})$ and $(T_{b_{i}},T_{a_{i}})$ are conjugate in $\operatorname{PMod}(\Sigma)$: we can find $f\in\operatorname{PMod}(\Sigma)$ with $f(a_{i})=b_{i}$ and $f(b_{i})=a_{i}$, so that $fT_{a_{i}}f^{-1}=T_{b_{i}}$ and $fT_{b_{i}}f^{-1}=T_{a_{i}}$. The operators $M_{i}N_{i}$ and $N_{i}M_{i}$ are thus conjugated by $\rho(f)\in\operatorname{GL}_{d}(\mathbb{C})$. We are done. ∎